# A Step Toward World Models: A Survey on Robotic Manipulation

### v1

@@ V-C Long-horizon Reasoning
Many robotic tasks require coherent long-horizon reasoning, where achieving the final objective depends on executing a temporally consistent sequence of actions over extended time scales. Existing methods are limited in long-horizon predictions [76, 4, 29, 13, 15]. For example, Ha et al. [4, 29, 13, 15] predefine temporal horizons to guide planning in their world models. In terms of video generation, existing methods still suffer from limited length (short-horizon future video) [65]. For example, Ko et al. [70] predicts a fixed number (eight) of future frames with U-Net based diffusion model [146]. Bruce et al. [20] can only memorize 16 frames and cannot produce consistent predictions. For autoregressive models, small prediction errors compound sequentially, leading to substantial inaccuracies in long-horizon forecasts.
V-C1 Closed-loop learning scheme
A line of work enabling long-term planning/predictions by learning through interaction with feedback and adjusting their behaviour accordingly [5, 22] . For example, Ebert et al. [30, 64] utilize image-to-image registration between predicted video frames and both the start and the goal images with the average length of the warping vectors as a cost function. The model would continue to retry until the task is completed. Du et al. [5] proposes a recursive planning framework comprising action proposal, video rollout generation, and evaluation. Vision–language models (VLMs) are used to propose potential next actions, while video generation models simulate multiple possible future rollouts. The resulting trajectories are then evaluated by the VLMs to select the optimal action. Du et al. [57] design a neural simulator that predicts future visuals, enabling policy models to interact within a consistent environment. A sparse memory mechanism is leveraged to further enhance the consistency over the time.
@@ V-C2 Subgoals
Pre-trained models possess a vast repository of commonsense and procedural knowledge that can be leveraged to decompose a high-level goal, often specified in natural language (e.g., ”make a cup of coffee”), into a logical sequence of concrete sub-goals or skills. Bu et al. [22] propose to promote long-horizon manipulation tasks by decomposing the goal into sub-goals and handling error accumulations by designing a real-time feedback mechanism. Yang et al. [10] leverage VLM to produce sub-goals and utilize coarse and fine video diffusion models to generate long-horizon videos. Chen et al. [150] utilizes an LLM to generate a multi-stage plan and design a LLM-based dense reward generator for sub-tasks, providing crucial guidance for long-horizon planning.
@@ V-C3 Hierarchical structures
Bu et al. [56] propose hierarchical world models with Adaptive Temporal Abstractions that separate the modeling of dynamics into high-level and low-level latent states. The low-level model captures fine-grained, short-term dynamics for immediate reactions, while the high-level model abstracts over longer temporal horizons to represent extended dependencies and long-term goals. By dynamically adapting the temporal granularity of the high-level latent states, the model can efficiently plan and predict over long horizons while maintaining accurate short-term predictions through the low-level module.
@@
Driess et al. [54] provide a goal image in addition to language instructions. Du et al. [5] propose to take advantage of long-horizon inference of VLMs and the low-level visual dynamic modelling ability of text-to-video models to handle long-horizon visual planning. A tree search over the space of possible video sequences to find proper long-horizon plans. Ren et al. [75] lean compact representations for the visual world that preserve the detailed temporal dynamics by means of causal encoder-decoder and quantization with a discrete codebook [148].
@@ V-D1 Data perspective
In conditional video synthesis, Du et al. [59] incorporates the observed image as additional context when denoising each frame. Specifically, it adapts a temporal super-resolution diffusion architecture by tiling the conditioned visual observation across all timesteps. Each intermediate noisy frame is concatenated with the observed image throughout sampling, providing a strong spatial anchor that enforces consistent environmental states across time. Ko et al. [70] concatenates the initial condition frame with all subsequent frames, providing a stable reference that preserves both the spatial layout and temporal evolution of the environment throughout the sequence. Zhen et al. [24] refine depth maps using normal integration to enhance spatial consistency. Optical flow is then calculated to ensure depth coherence across frames, maintaining consistent scene geometry over time.
@@ V-D2 Model perspective
Yang et al. [10] noted that in autoregressive predictions, standard spatiotemporal attention in video diffusion models degrades frame consistency due to limited long-range context. To address this, the temporal attention layers are replaced with 3D full attention layers, enabling computation of attention across all spatiotemporal tokens and better modeling of large motions. Additionally, the spatial attention layers are modified by reinjecting the VAE features of the first frame and computing cross-attention with the spatial tokens of the query features, further enhancing frame coherence.
@@ V-D3 Memory mechanism
is often used to enhance the spatiotemporal consistency. For example, Liao et al. [57] design a sparse memory mechanism to provide long-term historical context, improving spatiotemporal consistency and task relevance. More information can refer to Section V-G.
@@ V-E1 Data scaling
An intuitive and effective strategy to enhance generalization is to scale the diversity and volume of training data. For example, Cheang et al. [58] increase the number of pre-training videos from 0.8 million in [151] to 38 million. Assran et al. [77] expand the dataset from 2 million used by [152] to 22 million videos. Wang et al. [73] expand each of the 40 datasets by increasing trajectories from 10 up to $10^{6}$. Cheang et al. [153] train the model with web-scale vision-language data, human trajectory data and robot trajectory data. Kevin et al. [154] leverage diverse mobile manipulator data, diverse multi-environment non-mobile robot data, cross-embodiment laboratory data, high-level subtask prediction, and multi-modal web data. Cheang et al. [33, 58] investigate data augmentation strategies to enhance generalization. In [33], object rotation and roto-translation are applied. Cheang et al. [58] generate novel scenes by injecting objects using a diffusion model [155] and/or altering backgrounds with the Segment Anything Model (SAM) [107]. A video generation model [107] is subsequently employed to synthesize videos that preserve the original robot motions from the inpainted frames. Liao et al. [57] augment the dataset with a diverse set of failure cases, including erroneous executions, incomplete behaviors, and suboptimal control trajectories—collected from both human teleoperation and real-world robotic deployments. One problem of data scaling is that it is unlikely to collect all data for each tasks. At the same time, how to balance different data tasks is also challenging. Moreover, performance gains by scaling data is also limited for consistent performance improvements.
@@ V-E2 Use of pretrained models
Many methods aim to enhance generalization by leveraging the generative capabilities of video models. For example, Zhu et al. [23] combine video generation with geometric-aware learning to improve synthetic-to-real generalization across unseen viewpoints and support multiple downstream tasks. Zhen et al. [24] fine-tune a video generation model on RGB, depth, and normal videos to encode detailed shape, configuration, and temporal dynamics, enabling generalization to unseen scenes, objects, and cross-domain scenarios. The generalization capabilities of large language models, such as video-language models [86] and vision-language models [101], can be leveraged to enhance world models. By extracting high-level knowledge about the environment, these models facilitate more effective low-level dynamics modeling.
@@ V-E3 Instructions decomposing
Another generation issue comes from unseen instructions. To handle this, Zhou et al. [66] enhance the ability to unseen instructions by decomposing each spatial relation phrase into a set of compositional components with the pre-trained parser [156] and the rule-based approach. Detailed information can refer to Section V-B2a.
@@ V-E4 Invariant Representations
Generalization can be significantly improved by learning representations that are invariant to superficial or task-irrelevant changes in the environment. For example, Pang et al. [157] model learns to explicitly decompose visual observations into a view-invariant representation, which is used for the control policy, and a view-dependent representation. This decoupling makes the resulting policy robust to changes in camera viewpoint, a common source of failure in visuomotor control. Similarly, the Martinez et al. [158] framework learns a transferable communicative context between two agents, which enables zero-shot adaptation to entirely unseen sparse-reward environments by decoupling the representation learning from the control problem. Wu et al. [159] disentangle the modeling of context and dynamics by introducing a context encoder, enabling the model to capture shared knowledge for predictions.
@@ V-E5 Task-relevant information focused
Video data often contain irrelevant data to the actions such as background and robot arm, which would limited the generalization ability of the learned world models. To handle this, [69] propose to object-centric world models, which concentrated on object movements via the optical flow predictions that is independent of embodiment. Finn et al. [36] propose to explicitly model and predict motion that are relatively invariant to the object appearance, enabling long-range predictions and generalize to unseen objects.
@@ V-E6 Other strategies
Black et al. [68] use a pretrained image-editing model to generate subgoals from language commands and current observations, enabling low-level controllers to act and generalize to novel objects and scenarios. Self-supervised learning without task-specific rewards that can enhancing generalization abilities into different tasks [39].
@@ V-F Physics-informed Learning
Existing world models struggle to generate physically consistent videos because they lack an inherent understanding of physics, often producing unrealistic dynamics and implausible event sequences. Simply scaling up training data or model size is insufficient to capture the underlying physical laws [160]. To address this challenge, several approaches have been proposed. For example, Yang et al. [161] introduce a two-stage image-to-video generation framework that explicitly incorporates physics through vision- and language-informed physical priors. Team et al. [23] estimate depth and camera pose directly from videos, facilitating physics-informed learning and enabling world models to infer and predict physically consistent dynamics. Peper et al. [162] argue that advancing from physics-informed to physics-interpretable world models requires rethinking model design, and propose four guiding principles: organizing latent spaces by physical intent, encoding invariant and equivariant environmental representations, integrating multiple supervision signals, and partitioning generative outputs to improve both scalability and verifiability.
@@ V-G Memory
Memory mechanisms enable world models to store and retrieve relevant past information, supporting hidden-state disambiguation and long-horizon reasoning. For example, LeCun et al. [11] incorporate a memory module that maintains past, current, and predicted world states along with intrinsic costs, allowing retrieval of contextual information for reasoning and training. Huang et al. [8] propose a sparse contextual memory mechanism that preserves essential prior information throughout the generation process in a non-redundant manner, theoretically enabling the generation of sequences of arbitrary length. Zhou et al. [163] employ a 3D feature-map memory to maintain temporal consistency during sequence generation.
Memory efficiency Standard transformer blocks apply Multi-Head Self-Attention (MHA) to all tokens in the input token sequence, resulting in quadratic computation cost. Zhu et al. [43] leverage the memory-efficient spatial-temporal attention mechanism to reduce the computation cost. Liao et al. [57] randomly sampled parse memory frames from prior video history to augment temporal diversity to improve representational invariance, and use low-frame-rate video sequence for fine-tuning frames.
V-H Other Challenges
@@ V-H1 Video fidelity
To achieve high-fidelity video generation, several methods leverage powerful generative models. For instance, Ko et al. [70] employ an image diffusion model based on a U-Net with factorized spatiotemporal convolutions as the fundamental building block. Guo et al. [21] utilize the pre-trained variational autoencoder from Stable Diffusion [164]. Souvcek et al. [165] propose to make use of a variety of action and final state prompts.
@@ V-H2 Closed-loop Learning
Closed-loop learning enables agents to actively refine their internal world models by observing and responding to real-time feedback from the environment. This continuous perception–action cycle grounds learning in physical reality, enhances generalization, and allows adaptive correction—key properties for robust embodied intelligence. Driess et al. [54] update observations based on the actions executed, which are then fed into VLMs to enable the robot to correct or reorganize its plan in response to environmental changes and task progress. Bu et al. [22] design a feedback mechanism that is based on the element-wise discrepancy measure between current and goal state embeddings. Zhi et al. [69], estimate the location of the moving objects, depth prediction, 3D optical flow by input into GPT-4o to verify alignment with given instructions, enabling closed-loop planning.
V-H3 Sim-to-real gap
Huang et al. [8] propose combining the generative model with 4D Gaussian Splatting, forming a self-reinforcing data loop to reduce the sim-to-real gap.
V-H4 3D robotics world predicting
General-purpose video generation models neglect the substantial gap between their representation space and the three-dimensional, temporally interconnected robotics environment, thereby hindering accurate action policy prediction. For example, Wen et al. [166] focuses on 2D image prediction before action generation. To handle this, Huang et al. [8] propose Free Anchor Views (FAVs), a multi-view video representation offering flexible, task-adaptive perspectives to address challenges like motion ambiguity and environmental constraints.
V-H5 Fine-grained robot-object interaction
Robots are expected to perform precise manipulation, which requires world models to support fine-grained robot-object interactions. To achieve this, Zhu et al. [43] design a novel frame-level action-conditioning module to achieve precise action-frame alignment. He et al. [133] adopt two different pre-trained video generative models as the base models, introduces a minimalist yet powerful add-on action-conditioned module that improves frame-level action awareness while maintaining architectural flexibility.
V-H6 Multi-agent operation
Certain tasks necessitate coordinated operation among multiple robots to achieve successful completion. To this end, Zhang et al. [7] factorize the joint actions of different agents as a set of text prompt and leverage composable video diffusion models to learn world dynamics and make predictions. An agent-dependent loss is imposed to let the model focus on the related pixel, where the loss coefficient matrix is based on each agent’s reachable region reachable region.
V-H7 Reasoning
@@
Zhou et al. [60] enhance the reasoning and genrealization ability by incorporating context information and predicting dynamic regions, depth map, semantic knowledge by means of foundation models, e.g., DINOv2 [106] and SAM [107]. Ye et al. [115] introduce an Embodied Chain-of-Thought (CoT) as an intermediate reasoning representation, enabling more structured and interpretable decision-making in embodied tasks. Ye et al. [115] [79] generates a sub-goal image that represents the robot’s planned state in pixel space, and then conditions its action on both the current observation and the generated subgoal image.

### v2

@@ V-C Long-horizon Reasoning
Many robotic tasks require coherent long-horizon reasoning, where achieving the final objective depends on executing a temporally consistent sequence of actions over extended time scales. Existing methods are limited in long-horizon predictions [nair2022learning, ha2018world, hafner2019learning, hafner2021mastering, hafner2023mastering]. For example, Ha et al. [ha2018world, hafner2019learning, hafner2021mastering, hafner2023mastering] predefine temporal horizons to guide planning in their world models. In terms of video generation, existing methods still suffer from limited length (short-horizon future video) [gao2024flip]. For example, Ko et al. [ko2024learning] predicts a fixed number (eight) of future frames with U-Net based diffusion model [dhariwal2021diffusion]. Bruce et al. [bruce2024genie] can only memorize 16 frames and cannot produce consistent predictions. For autoregressive models, small prediction errors compound sequentially, leading to substantial inaccuracies in long-horizon forecasts.
V-C1 Closed-loop Learning
A line of work enabling long-term planning/predictions by learning through interaction with feedback and adjusting their behaviour accordingly [du2023video, bu2024closed] . For example, Ebert et al. [ebert2018robustness, ebert2018visual] utilize image-to-image registration between predicted video frames and both the start and the goal images with the average length of the warping vectors as a cost function. The model would continue to retry until the task is completed. Du et al. [du2023video] proposes a recursive planning framework comprising action proposal, video rollout generation, and evaluation. Vision–language models (VLMs) are used to propose potential next actions, while video generation models simulate multiple possible future rollouts. The resulting trajectories are then evaluated by the VLMs to select the optimal action. Du et al. [liao2025genie] design a neural simulator that predicts future visuals, enabling policy models to interact within a consistent environment. A sparse memory mechanism is leveraged to further enhance the consistency over the time.
@@ V-C2 Subgoals
Pre-trained models possess a vast repository of commonsense and procedural knowledge that can be leveraged to decompose a high-level goal, often specified in natural language (e.g., ”make a cup of coffee”), into a logical sequence of concrete sub-goals or skills. Bu et al. [bu2024closed] propose to promote long-horizon manipulation tasks by decomposing the goal into sub-goals and handling error accumulations by designing a real-time feedback mechanism. Yang et al. [yang2025roboenvision] leverage VLM to produce sub-goals and utilize coarse and fine video diffusion models to generate long-horizon videos. Chen et al. [chen2025robohorizon] utilizes an LLM to generate a multi-stage plan and design a LLM-based dense reward generator for sub-tasks, providing crucial guidance for long-horizon planning.
@@ V-C3 Hierarchical Structures
Bu et al. [gumbsch2023learning] propose hierarchical world models with Adaptive Temporal Abstractions that separate the modeling of dynamics into high-level and low-level latent states. The low-level model captures fine-grained, short-term dynamics for immediate reactions, while the high-level model abstracts over longer temporal horizons to represent extended dependencies and long-term goals. By dynamically adapting the temporal granularity of the high-level latent states, the model can efficiently plan and predict over long horizons while maintaining accurate short-term predictions through the low-level module.
@@
Driess et al. [driess2023palm] provide a goal image in addition to language instructions. Du et al. [du2023video] propose to take advantage of long-horizon inference of VLMs and the low-level visual dynamic modelling ability of text-to-video models to handle long-horizon visual planning. A tree search over the space of possible video sequences to find proper long-horizon plans. Ren et al. [ren2025videoworld] lean compact representations for the visual world that preserve the detailed temporal dynamics by means of causal encoder-decoder and quantization with a discrete codebook [mentzer2024finite].
@@ V-D1 Data Perspective
In conditional video synthesis, Du et al. [du2023learning] incorporate the observed image as additional context when denoising each frame. Specifically, it adapts a temporal super-resolution diffusion architecture by tiling the conditioned visual observation across all timesteps. Each intermediate noisy frame is concatenated with the observed image throughout sampling, providing a strong spatial anchor that enforces consistent environmental states across time. Ko et al. [ko2024learning] concatenates the initial condition frame with all subsequent frames, providing a stable reference that preserves both the spatial layout and temporal evolution of the environment throughout the sequence. Zhen et al. [zhen2025tesseract] refine depth maps using normal integration to enhance spatial consistency. Optical flow is then calculated to ensure depth coherence across frames, maintaining consistent scene geometry over time.
@@ V-D2 Model Perspective
Yang et al. [yang2025roboenvision] noted that in autoregressive predictions, standard spatiotemporal attention in video diffusion models degrades frame consistency due to limited long-range context. To address this, the temporal attention layers are replaced with 3D full attention layers, enabling computation of attention across all spatiotemporal tokens and better modeling of large motions. Additionally, the spatial attention layers are modified by reinjecting the VAE features of the first frame and computing cross-attention with the spatial tokens of the query features, further enhancing frame coherence.
@@ V-D3 Memory Mechanism
Memory mechanisms preserve historical information, enabling the coherent evolution of spatial and temporal patterns over time. For example, Liao et al. [liao2025genie] design a sparse memory mechanism to provide long-term historical context, improving spatiotemporal consistency and task relevance. More information can refer to Section V-G.
@@ V-E1 Data Scaling
An intuitive and effective strategy to enhance generalization is to scale the diversity and volume of training data. For example, Cheang et al. [cheang2024gr] increase the number of pre-training videos from 0.8 million in [wu2024unleashing] to 38 million. Assran et al. [assran2025v] expand the dataset from 2 million used by [bardes2024revisiting] to 22 million videos. Wang et al. [wang2025learning] expand each of the 40 datasets by increasing trajectories from 10 up to $10^{6}$. Cheang et al. [cheang2025gr] train the model with web-scale vision-language data, human trajectory data and robot trajectory data. Kevin et al. [intelligence2025pi_] leverage diverse mobile manipulator data, diverse multi-environment non-mobile robot data, cross-embodiment laboratory data, high-level subtask prediction, and multi-modal web data. Cheang et al. [barcellona2025dream, cheang2024gr] investigate data augmentation strategies to enhance generalization. In [barcellona2025dream], object rotation and roto-translation are applied. Cheang et al. [cheang2024gr] generate novel scenes by injecting objects using a diffusion model [ho2020denoising] and/or altering backgrounds with the Segment Anything Model (SAM) [kirillov2023segment]. A video generation model [kirillov2023segment] is subsequently employed to synthesize videos that preserve the original robot motions from the inpainted frames. Liao et al. [liao2025genie] augment the dataset with a diverse set of failure cases, including erroneous executions, incomplete behaviors, and suboptimal control trajectories—collected from both human teleoperation and real-world robotic deployments. One problem of data scaling is that it is unlikely to collect all data for each tasks. At the same time, how to balance different data tasks is also challenging. Moreover, performance gains by scaling data is also limited for consistent performance improvements.
@@ V-E2 Use of Pretrained Models
Many methods aim to enhance generalization by leveraging the generative capabilities of video models. For example, Zhu et al. [team2025aether] combine video generation with geometric-aware learning to improve synthetic-to-real generalization across unseen viewpoints and support multiple downstream tasks. Zhen et al. [zhen2025tesseract] fine-tune a video generation model on RGB, depth, and normal videos to encode detailed shape, configuration, and temporal dynamics, enabling generalization to unseen scenes, objects, and cross-domain scenarios. The generalization capabilities of large language models, such as video-language models [wang2025founder] and vision-language models [mazzaglia2024genrl], can be leveraged to enhance world models. By extracting high-level knowledge about the environment, these models facilitate more effective low-level dynamics modeling.
@@ V-E3 Instructions Decomposing
Another generation issue comes from unseen instructions. To handle this, Zhou et al. [zhou2024robodreamer] enhance the ability to unseen instructions by decomposing each spatial relation phrase into a set of compositional components with the pre-trained parser [kitaev2019multilingual] and the rule-based approach. Detailed information can refer to Section V-B2a.
@@ V-E4 Invariant Representations
Generalization can be significantly improved by learning invariant representations to superficial or task-irrelevant changes in the environment. For example, Pang et al. [pang2025reviwo] model learns to explicitly decompose visual observations into a view-invariant representation, which is used for the control policy, and a view-dependent representation. This decoupling makes the resulting policy robust to changes in camera viewpoint, a common source of failure in visuomotor control. Similarly, the Martinez et al. [martinez2025coral] framework learns a transferable communicative context between two agents, which enables zero-shot adaptation to entirely unseen sparse-reward environments by decoupling the representation learning from the control problem. Wu et al. [wu2023pre] disentangle the modeling of context and dynamics by introducing a context encoder, enabling the model to capture shared knowledge for predictions.
@@ V-E5 Task-relevant Information Focused
Video data often contain irrelevant data to the actions such as background and robot arm, which would limited the generalization ability of the learned world models. To handle this, [zhi20253dflowaction] propose to object-centric world models, which concentrated on object movements via the optical flow predictions that is independent of embodiment. Finn et al. [finn2016unsupervised] propose to explicitly model and predict motion that are relatively invariant to the object appearance, enabling long-range predictions and generalize to unseen objects.
@@ V-E6 Other Strategies
Black et al. [black2024zero] use a pretrained image-editing model to generate subgoals from language commands and current observations, enabling low-level controllers to act and generalize to novel objects and scenarios. Self-supervised learning without task-specific rewards that can enhancing generalization abilities into different tasks [sekar2020planning].
@@ V-F Physics-informed Learning
Existing world models struggle to generate physically consistent videos because they lack an inherent understanding of physics, often producing unrealistic dynamics and implausible event sequences. Simply scaling up training data or model size is insufficient to capture the underlying physical laws [kang2025far]. To address this challenge, several approaches have been proposed. For example, Yang et al. [yang2025vlipp] introduce a two-stage image-to-video generation framework that explicitly incorporates physics through vision- and language-informed physical priors. Team et al. [team2025aether] estimate depth and camera pose directly from videos, facilitating physics-informed learning and enabling world models to infer and predict physically consistent dynamics. Peper et al. [peper2025four] argue that advancing from physics-informed to physics-interpretable world models requires rethinking model design, and propose four guiding principles: organizing latent spaces by physical intent, encoding invariant and equivariant environmental representations, integrating multiple supervision signals, and partitioning generative outputs to improve both scalability and verifiability.
@@ V-G Memory
Memory mechanisms enable world models to store and retrieve relevant past information, supporting hidden-state disambiguation and long-horizon reasoning. For example, LeCun et al. [lecun2022path] incorporate a memory module that maintains past, current, and predicted world states along with intrinsic costs, allowing retrieval of contextual information for reasoning and training. Huang et al. [huang2025enerverse] propose a sparse contextual memory mechanism that preserves essential prior information throughout the generation process in a non-redundant manner, theoretically enabling the generation of sequences of arbitrary length. Zhou et al. [zhou2025learning] employ a 3D feature-map memory to maintain temporal consistency during sequence generation.
Memory Efficiency. Standard transformer blocks apply multi-head self-attention to all tokens in the input token sequence, resulting in quadratic computation cost. Zhu et al. [zhu2025irasim] leverage the memory-efficient spatial-temporal attention mechanism to reduce the computation cost. Liao et al. [liao2025genie] randomly sampled parse memory frames from prior video history to augment temporal diversity to improve representational invariance, and use low-frame-rate video sequence for fine-tuning frames.
V-H Other Techniques and Challenges
@@ V-H1 Video Fidelity
To achieve high-fidelity video generation, several methods leverage powerful generative models. For instance, Ko et al. [ko2024learning] employ an image diffusion model based on a U-Net with factorized spatiotemporal convolutions as the fundamental building block. Guo et al. [guo2025flowdreamer] utilize the pre-trained variational autoencoder from Stable Diffusion [rombach2022high]. Souvcek et al. [souvcek2024genhowto] propose to make use of a variety of action and final state prompts.
@@ V-H2 Closed-loop Learning
Closed-loop learning enables agents to actively refine their internal world models by observing and responding to real-time feedback from the environment. This continuous perception–action cycle grounds learning in physical reality, enhances generalization, and allows adaptive correction. Driess et al. [driess2023palm] update observations based on the actions executed, which are then fed into VLMs to enable the robot to correct or reorganize its plan in response to environmental changes and task progress. Bu et al. [bu2024closed] design a feedback mechanism that is based on the element-wise discrepancy measure between current and goal state embeddings. Zhi et al. [zhi20253dflowaction], estimate the location of the moving objects, depth prediction, 3D optical flow by input into GPT-4o to verify alignment with given instructions, enabling closed-loop planning.
V-H3 Reasoning
@@
Reasoning enables a robot to interpret sensory input and translate it into purposeful actions rather than reactive responses. Zhou et al. [zhang2025dreamvla] enhance the reasoning and genrealization ability by incorporating context information and predicting dynamic regions, depth map, semantic knowledge by means of foundation models, e.g., DINOv2 [oquab2024dinov2] and SAM [kirillov2023segment]. Ye et al. [Ye2025GigaBrain] introduce an Embodied Chain-of-Thought as an intermediate reasoning representation, enabling more structured and interpretable decision-making in embodied tasks. Ye et al. [Ye2025GigaBrain] [zhao2025cot] generates a sub-goal image that represents the robot’s planned state in pixel space, and then conditions its action on both the current observation and the generated subgoal image.
V-H4 Sim-to-real Gap
The substantial gap between synthetic and real-world data limits the transferability and real-world performance of robotic systems. To handle this, Huang et al. [huang2025enerverse] propose combining the generative model with 4D Gaussian Splatting, forming a self-reinforcing data loop to reduce the sim-to-real gap.
V-H5 3D Robotics World Prediction
General-purpose video generation models neglect the substantial gap between their representation space and the three-dimensional, temporally interconnected robotics environment, thereby hindering accurate action policy prediction. For example, Wen et al. [wen2024vidman] focuses on 2D image prediction before action generation. To handle this, Huang et al. [huang2025enerverse] propose Free Anchor Views, a multi-view video representation offering flexible, task-adaptive perspectives to address challenges like motion ambiguity and environmental constraints.
V-H6 Fine-grained Robot-object Interaction
Robots are expected to perform precise manipulation, which requires world models to support fine-grained robot-object interactions. To achieve this, Zhu et al. [zhu2025irasim] design a novel frame-level action-conditioning module to achieve precise action-frame alignment. He et al. [he2025pre] adopt two different pre-trained video generative models as the base models, and introduce a minimalist yet powerful add-on action-conditioned module that improves frame-level action awareness while maintaining architectural flexibility.
V-H7 Multi-agent Operation
Certain tasks necessitate coordinated operation among multiple robots to achieve successful completion. To this end, Zhang et al. [zhang2025combo] factorize the joint actions of different agents as a set of text prompt and leverage composable video diffusion models to learn world dynamics and make predictions. An agent-dependent loss is imposed to let the model focus on the related pixel, where the loss coefficient matrix is based on each agent’s reachable region.